A plane flies 350 mi with the wind in the same time that it can fly 310 against the wind. The plane's speed is 165 mph in still air. Find the speed of the wind.
Answer by mananth(16949)

let speed of wind be x mph
..
Distance traveled against wind =310
speed against wind = 165-x
Time taken = 310/165-x
..
Distance traveled with wind =350 miles.
speed with wind = x+165
time taken = 350/x+165
..
Both ways the time is same
310/165-x= 350/x+165
310(165-x)=350(165+x)
51150-310x = 57750+350x
660x=6600
x=10mph the speed of wind